OpenStudy (anonymous):

The measure of an interior angle of a regular polygon is 165.6 . How many sides make up the regular polygon?

OpenStudy (anonymous):

Do you know what the sum of the angles of the polygon are? ;-) i.e.: Triangle = 180 , any Quadrilateral = 360, etc.

OpenStudy (anonymous):

i think its 165.6 but im not entirely sure

OpenStudy (anonymous):

There are an infinite number of regular polygons (and as you get towards infinity it's easier to say it's just a circle)

OpenStudy (anonymous):

Did you mean? "The SUM of the interior angles of a regular polygon is 165.6\(^o\)?"

OpenStudy (anonymous):

Thisssssss is a trap question mate!

OpenStudy (anonymous):

all it says is the measure. no sum

OpenStudy (anonymous):

Badly worded question then (I take it, it was a given question, not one of your own), but no matter, here's what I think you need: For any REGULAR polygon with "n" sides: \(\huge\cdot \) Total sum of interior angles = (n-2) \(\cdot\) (180° or \(\large\frac{\pi}{2}\)) \(\huge\cdot \) Each angle = \(\large \frac{(n-2) \cdot (180° or \large\frac{\pi}{2})}{n}\)

OpenStudy (anonymous):

Write it down sssssssssssir :-3

OpenStudy (anonymous):

Ack! That should be Pi not Pi/2

OpenStudy (anonymous):

180\(^o\) = \(\pi\) radians, not \(\frac{\pi}{2}\)

OpenStudy (anonymous):

My bad I made a typo for that, but otherwise the formulas above are golden. You're probably using degrees anyways I would guess.